# Spin waves in the bilayer van der Waals magnet CrSBr

### v1

@@ II.1 Hamiltonian
Parameter values for CrSBr vary across the literature [30, 31]. Table I summarizes the numerical values used in this work. We adopt the parameters for the monolayer from density functional theory (DFT) [22], where the exchange constants are multiplied by a factor 1/2 and the sign of the anisotropy constants is changed to remain consistent with our choice of Hamiltonian. The calculated interlayer exchange coupling constants [22] are three orders of magnitude smaller than the intralayer coupling. However, for the interlayer exchange coupling, we prefer the experimental value $\mathcal{J}_{\perp}=-15.5\mu eV$ [32] with a modulus that is an order of magnitude larger than the estimate from DFT. Both the spin-orbit coupling (SOC) and shape anisotropy originating from dipole-dipole interactions contribute to the total magnetic anisotropy of monolayer CrSBr [22]. In the ground state, the second term dominates. Finally, the saturation magnetization of the CrSBr monolayer is $M_{s}=0.48$ T [8, 33].
Table 1: Numerical values of the in-plane lattice constants a and b, the exchange constants $\mathcal{J}_{i}$, and the anisotropy constants $\mathcal{D}_{i}$ [22].
@@
$\mathcal{J}_{1}$ 1.61 $\mathrm{meV}$
$\mathcal{J}_{2}$ 2.06 $\mathrm{meV}$
$\mathcal{J}_{3}$ 1.39 $\mathrm{meV}$
@@
$\alpha_{1}=4(J_{1}+J_{3}),\ \alpha_{2}=8J_{2},$ (9)
$\beta_{1}=4[J_{1}\cos(k_{x}a)+J_{3}\cos(k_{z}b)],$
$\beta_{2}=4J_{2}\left[\cos\left(\frac{1}{2}k_{x}a+\frac{1}{2}k_{z}b\right)+\cos\left(\frac{1}{2}k_{x}a-\frac{1}{2}k_{z}b\right)\right].$

### v2

@@ II.1 Hamiltonian
Parameter values for CrSBr vary across the literature [30, 31, 26, 17]. Table I summarizes the numerical values used in this work. We adopt the parameters for the monolayer from density functional theory (DFT). The intralayer exchange coupling constants are taken from Ref. [13] and the magnetic anisotropy constants from Ref. [22], where the sign of the anisotropy constants is changed to remain consistent with our choice of Hamiltonian. For the interlayer exchange coupling, we prefer the experimental value $\mathcal{J}_{\perp}=-15.5\mu eV$ [32] with a modulus that is an order of magnitude larger than the estimate from DFT [22]. Both the spin-orbit coupling (SOC) and shape anisotropy originating from dipole-dipole interactions contribute to the total magnetic anisotropy of monolayer CrSBr [22]. In the ground state, the second term dominates. Finally, the saturation magnetization of the CrSBr monolayer is $M_{s}=0.48$ T [8, 33].
Table 1: Numerical values of the in-plane lattice constants a and b, the exchange constants $\mathcal{J}_{i}$ [13], and the anisotropy constants $\mathcal{D}_{i}$ [22].
@@
$\mathcal{J}_{1}$ 3.54 $\mathrm{meV}$
$\mathcal{J}_{2}$ 3.08 $\mathrm{meV}$
$\mathcal{J}_{3}$ 4.15 $\mathrm{meV}$
@@
$\alpha_{1}=2(J_{1}+J_{3}),\ \alpha_{2}=4J_{2},$ (9)
$\beta_{1}=2[J_{1}\cos(k_{x}a)+J_{3}\cos(k_{z}b)],$
$\beta_{2}=2J_{2}\left[\cos\left(\frac{1}{2}k_{x}a+\frac{1}{2}k_{z}b\right)+\cos\left(\frac{1}{2}k_{x}a-\frac{1}{2}k_{z}b\right)\right].$